16. GEOMETRY A triangle has an area of 3x2+ 9x-54 square centimeters. f the height of the triangle is x+6 centimeters, find the length of the base.
Answer #1

Given that the area of triangle is

(3x2 + 9x - 54) cm2

Also the height of triangle is

(x+ 6) cm

We know that

Area of a triangle = (1/2) *length of Base * Height

length of Base= 2* Area of triangle / Height

=2 (3x2 + 9x - 54) /   (x+ 6)

= 6 (x2 + 3x - 18) /   (x+ 6)

= 6(x+6)(x-3) / (x+6)

Length of base = 6 (x-3) cm